# Inv-Entropy: A Fully Probabilistic Framework for Uncertainty Quantification in Language Models

### v1

@@ D.1 Experimental setting
D.2 Additional experimental results
@@
Table 9 presents the results obtained using LLaMA. Inv-Entropy continues to demonstrate strong performance, achieving competitive or state-of-the-art results across most datasets and evaluation metrics.
Dataset Metric $S{=}1$ $S{=}5$ $S{=}10$ $S{=}30$ $S{=}50$ $S{=}100$
TriviaQA AUROC 0.743$\pm$0.061 0.781$\pm$0.057 0.785$\pm$0.056 0.788$\pm$0.054 0.780$\pm$0.054 0.788$\pm$0.054
PRR 0.840$\pm$0.054 0.881$\pm$0.043 0.882$\pm$0.043 0.885$\pm$0.044 0.882$\pm$0.044 0.885$\pm$0.044
Brier 0.138$\pm$0.021 0.127$\pm$0.020 0.125$\pm$0.020 0.128$\pm$0.020 0.131$\pm$0.021 0.128$\pm$0.020
SciQ AUROC 0.724$\pm$0.052 0.733$\pm$0.050 0.740$\pm$0.049 0.740$\pm$0.050 0.739$\pm$0.050 0.743$\pm$0.049
PRR 0.821$\pm$0.052 0.840$\pm$0.046 0.844$\pm$0.045 0.843$\pm$0.042 0.842$\pm$0.046 0.845$\pm$0.046
Brier 0.163$\pm$0.016 0.160$\pm$0.015 0.159$\pm$0.015 0.157$\pm$0.018 0.160$\pm$0.015 0.159$\pm$0.015
NQ AUROC 0.659$\pm$0.063 0.686$\pm$0.060 0.699$\pm$0.057 0.703$\pm$0.060 0.702$\pm$0.058 0.705$\pm$0.059
PRR 0.709$\pm$0.068 0.730$\pm$0.071 0.760$\pm$0.065 0.764$\pm$0.064 0.764$\pm$0.063 0.766$\pm$0.064
Brier 0.199$\pm$0.019 0.194$\pm$0.020 0.184$\pm$0.019 0.182$\pm$0.020 0.183$\pm$0.020 0.182$\pm$0.021
MMLU AUROC 0.604$\pm$0.059 0.762$\pm$0.042 0.777$\pm$0.042 0.780$\pm$0.041 0.790$\pm$0.041 0.789$\pm$0.040
PRR 0.743$\pm$0.064 0.863$\pm$0.046 0.897$\pm$0.031 0.898$\pm$0.030 0.905$\pm$0.028 0.902$\pm$0.030
Brier 0.188$\pm$0.021 0.152$\pm$0.017 0.148$\pm$0.017 0.147$\pm$0.017 0.142$\pm$0.017 0.143$\pm$0.017
GSM8K AUROC 0.674$\pm$0.054 0.684$\pm$0.056 0.685$\pm$0.054 0.695$\pm$0.051 0.690$\pm$0.055 0.695$\pm$0.055
PRR 0.487$\pm$0.090 0.530$\pm$0.096 0.527$\pm$0.095 0.521$\pm$0.094 0.520$\pm$0.094 0.527$\pm$0.095
Brier 0.159$\pm$0.021 0.151$\pm$0.022 0.150$\pm$0.021 0.152$\pm$0.020 0.152$\pm$0.020 0.146$\pm$0.021
@@
Inv-Entropy AUROC 0.812$\pm$0.044 0.815$\pm$0.044 0.820$\pm$0.044 0.794$\pm$0.044 0.801$\pm$0.043 0.806$\pm$0.042
PRR 0.916$\pm$0.028 0.915$\pm$0.029 0.915$\pm$0.029 0.888$\pm$0.038 0.888$\pm$0.039 0.892$\pm$0.038
Brier 0.128$\pm$0.020 0.121$\pm$0.020 0.123$\pm$0.020 0.143$\pm$0.019 0.140$\pm$0.019 0.136$\pm$0.019
NI-Entropy AUROC 0.702$\pm$0.072 0.631$\pm$0.081 0.617$\pm$0.083 0.729$\pm$0.056 0.731$\pm$0.058 0.746$\pm$0.055
PRR 0.813$\pm$0.054 0.765$\pm$0.060 0.751$\pm$0.059 0.805$\pm$0.055 0.805$\pm$0.056 0.809$\pm$0.054
Brier 0.133$\pm$0.024 0.152$\pm$0.023 0.150$\pm$0.024 0.147$\pm$0.021 0.146$\pm$0.021 0.144$\pm$0.021
WD-px-py AUROC 0.762$\pm$0.058 0.771$\pm$0.058 0.772$\pm$0.057 0.684$\pm$0.052 0.688$\pm$0.058 0.685$\pm$0.057
PRR 0.866$\pm$0.046 0.875$\pm$0.044 0.876$\pm$0.044 0.825$\pm$0.050 0.818$\pm$0.054 0.819$\pm$0.051
Brier 0.134$\pm$0.021 0.131$\pm$0.021 0.130$\pm$0.021 0.173$\pm$0.017 0.169$\pm$0.020 0.172$\pm$0.019
MAX-py-x AUROC 0.782$\pm$0.047 0.794$\pm$0.045 0.804$\pm$0.044 0.754$\pm$0.049 0.760$\pm$0.051 0.757$\pm$0.049
PRR 0.904$\pm$0.028 0.910$\pm$0.027 0.915$\pm$0.026 0.864$\pm$0.047 0.862$\pm$0.049 0.861$\pm$0.048
Brier 0.134$\pm$0.019 0.130$\pm$0.019 0.124$\pm$0.019 0.154$\pm$0.018 0.155$\pm$0.019 0.154$\pm$0.017
@@
TriviaQA AUROC SBERT-small 0.792$\pm$0.051 0.666$\pm$0.072 0.833$\pm$0.050 0.835$\pm$0.046
SBERT-large 0.772$\pm$0.057 0.617$\pm$0.083 0.804$\pm$0.044 0.816$\pm$0.044
DeBERTa 0.788$\pm$0.054 0.786$\pm$0.057 0.518$\pm$0.060 0.723$\pm$0.054
PRR SBERT-small 0.899$\pm$0.037 0.803$\pm$0.053 0.920$\pm$0.029 0.920$\pm$0.029
SBERT-large 0.876$\pm$0.044 0.751$\pm$0.059 0.915$\pm$0.026 0.915$\pm$0.030
DeBERTa 0.885$\pm$0.044 0.883$\pm$0.043 0.763$\pm$0.051 0.875$\pm$0.038
Brier SBERT-small 0.127$\pm$0.021 0.151$\pm$0.022 0.106$\pm$0.021 0.114$\pm$0.021
SBERT-large 0.130$\pm$0.021 0.150$\pm$0.024 0.124$\pm$0.019 0.124$\pm$0.020
DeBERTa 0.128$\pm$0.020 0.124$\pm$0.020 0.184$\pm$0.019 0.148$\pm$0.019
SciQ AUROC SBERT-small 0.774$\pm$0.049 0.750$\pm$0.054 0.655$\pm$0.058 0.742$\pm$0.050
SBERT-large 0.806$\pm$0.042 0.746$\pm$0.055 0.685$\pm$0.057 0.757$\pm$0.049
DeBERTa 0.740$\pm$0.050 0.681$\pm$0.056 0.303$\pm$0.060 0.674$\pm$0.054
PRR SBERT-small 0.874$\pm$0.044 0.820$\pm$0.059 0.796$\pm$0.054 0.852$\pm$0.049
SBERT-large 0.892$\pm$0.038 0.809$\pm$0.054 0.819$\pm$0.051 0.861$\pm$0.048
DeBERTa 0.843$\pm$0.042 0.781$\pm$0.053 0.587$\pm$0.056 0.821$\pm$0.048
Brier SBERT-small 0.147$\pm$0.021 0.150$\pm$0.021 0.181$\pm$0.019 0.160$\pm$0.018
SBERT-large 0.136$\pm$0.019 0.144$\pm$0.021 0.172$\pm$0.019 0.154$\pm$0.017
DeBERTa 0.157$\pm$0.018 0.164$\pm$0.017 0.212$\pm$0.016 0.177$\pm$0.017
MMLU AUROC SBERT-small 0.689$\pm$0.060 0.576$\pm$0.063 0.723$\pm$0.049 0.704$\pm$0.056
SBERT-large 0.634$\pm$0.058 0.532$\pm$0.057 0.670$\pm$0.055 0.676$\pm$0.057
DeBERTa 0.780$\pm$0.041 0.710$\pm$0.052 0.573$\pm$0.061 0.585$\pm$0.059
PRR SBERT-small 0.812$\pm$0.064 0.719$\pm$0.069 0.869$\pm$0.035 0.832$\pm$0.057
SBERT-large 0.790$\pm$0.059 0.695$\pm$0.068 0.834$\pm$0.049 0.820$\pm$0.056
DeBERTa 0.898$\pm$0.030 0.823$\pm$0.055 0.777$\pm$0.054 0.749$\pm$0.062
Brier SBERT-small 0.170$\pm$0.019 0.192$\pm$0.020 0.163$\pm$0.018 0.165$\pm$0.019
SBERT-large 0.185$\pm$0.020 0.200$\pm$0.020 0.177$\pm$0.019 0.173$\pm$0.020
DeBERTa 0.147$\pm$0.017 0.168$\pm$0.021 0.188$\pm$0.018 0.189$\pm$0.020
@@
AUROC($\uparrow$) Semantic Entropy 0.579$\pm$0.044 0.679$\pm$0.052 0.700$\pm$0.062 0.732$\pm$0.051
VU 0.695$\pm$0.060 0.625$\pm$0.058 0.629$\pm$0.063 0.578$\pm$0.067
P(True) 0.604$\pm$0.050 0.592$\pm$0.046 0.571$\pm$0.043 0.583$\pm$0.038
LexSim 0.649$\pm$0.055 0.715$\pm$0.051 0.775$\pm$0.055 0.792$\pm$0.051
DegMat 0.734$\pm$0.056 0.688$\pm$0.067 0.794$\pm$0.049 0.730$\pm$0.059
LUQ 0.637$\pm$0.067 0.712$\pm$0.059 0.748$\pm$0.063 0.817$\pm$0.044
KLE 0.333$\pm$0.054 0.327$\pm$0.059 0.221$\pm$0.056 0.155$\pm$0.042
Inv-Entropy 0.870$\pm$0.044 0.795$\pm$0.051 0.827$\pm$0.043 0.810$\pm$0.042
NI-Entropy 0.762$\pm$0.063 0.666$\pm$0.073 0.634$\pm$0.083 0.673$\pm$0.055
WD-px-py 0.829$\pm$0.044 0.827$\pm$0.051 0.816$\pm$0.054 0.829$\pm$0.039
MAX-py-x 0.854$\pm$0.043 0.832$\pm$0.046 0.856$\pm$0.040 0.846$\pm$0.041
PRR($\uparrow$) Semantic Entropy 0.787$\pm$0.044 0.803$\pm$0.046 0.834$\pm$0.043 0.811$\pm$0.045
VU 0.836$\pm$0.044 0.791$\pm$0.041 0.798$\pm$0.053 0.727$\pm$0.051
P(True) 0.797$\pm$0.042 0.760$\pm$0.044 0.776$\pm$0.042 0.726$\pm$0.046
LexSim 0.810$\pm$0.045 0.824$\pm$0.040 0.877$\pm$0.043 0.854$\pm$0.043
DegMat 0.882$\pm$0.041 0.816$\pm$0.053 0.895$\pm$0.037 0.812$\pm$0.054
LUQ 0.854$\pm$0.043 0.856$\pm$0.042 0.874$\pm$0.048 0.893$\pm$0.039
KLE 0.704$\pm$0.048 0.646$\pm$0.050 0.623$\pm$0.051 0.516$\pm$0.049
Inv-Entropy 0.939$\pm$0.028 0.900$\pm$0.037 0.936$\pm$0.022 0.903$\pm$0.037
NI-Entropy 0.862$\pm$0.043 0.799$\pm$0.054 0.777$\pm$0.060 0.765$\pm$0.055
WD-px-py 0.938$\pm$0.021 0.918$\pm$0.029 0.912$\pm$0.043 0.923$\pm$0.025
MAX-py-x 0.932$\pm$0.029 0.920$\pm$0.030 0.946$\pm$0.021 0.913$\pm$0.039
Brier($\downarrow$) Semantic Entropy 0.166$\pm$0.023 0.150$\pm$0.026 0.141$\pm$0.023 0.156$\pm$0.023
VU 0.160$\pm$0.022 0.178$\pm$0.017 0.175$\pm$0.023 0.203$\pm$0.016
P(True) 0.172$\pm$0.022 0.188$\pm$0.020 0.179$\pm$0.021 0.198$\pm$0.020
LexSim 0.151$\pm$0.024 0.146$\pm$0.021 0.128$\pm$0.024 0.127$\pm$0.021
DegMat 0.140$\pm$0.021 0.149$\pm$0.022 0.115$\pm$0.018 0.145$\pm$0.023
LUQ 0.148$\pm$0.020 0.142$\pm$0.021 0.121$\pm$0.023 0.121$\pm$0.018
KLE 0.188$\pm$0.021 0.199$\pm$0.020 0.192$\pm$0.021 0.218$\pm$0.014
Inv-Entropy 0.085$\pm$0.019 0.127$\pm$0.021 0.117$\pm$0.019 0.132$\pm$0.017
NI-Entropy 0.104$\pm$0.021 0.151$\pm$0.022 0.142$\pm$0.024 0.164$\pm$0.018
WD-px-py 0.115$\pm$0.018 0.102$\pm$0.022 0.117$\pm$0.021 0.121$\pm$0.017
MAX-py-x 0.103$\pm$0.019 0.116$\pm$0.021 0.108$\pm$0.019 0.114$\pm$0.016

### v2

@@ D.1 Experimental setting
D.2 Computational cost comparison
For non-locally hosted LLMs, uncertainty quantification methods generally involve two stages: obtaining responses from the model via API calls and computing the uncertainty scores based on those responses. This applies to both our method and all baselines; the only exception is Verbalized Uncertainty (VU), which directly returns a score from the API without requiring post-processing. As discussed in the paper, any method that relies on perturbations and/or replications incurs additional computational cost. However, a major advantage is that responses for perturbed inputs can be generated in parallel, making the process more scalable in practice.
With this in mind, we divide the computational cost into two components:
  • (A)
    Computation time after responses are collected, and
  • (B)
    Total cost, which includes (i) along with API call time and perturbation generation.
Regarding API cost:
Our method introduces $n$ perturbations per question and generates $r$ response replications for each version, resulting in a total of $(n+1)\times r$ API calls per input. Thus, the cost scales linearly with both the number of perturbations and replications. That said, while our method introduces perturbations, it requires far fewer replications than several existing baselines. For instance, Semantic Entropy relies heavily on large $r$ values (with $n=0$); prior work recommends $r>10$ for stable performance [farquhar2024detecting] In contrast, we demonstrate that even without replication ($r=1$), our method remains effective. This is evidenced by the strong performance of NR-Inv-Entropy, which uses perturbation alone and still consistently outperforms many baselines.
Regarding computation after responses are collected:
Below, we present the compute results in Table 4 without parallelization. The reported numbers are averaged over all sampled questions in TriviaQA. As shown, our method is highly efficient in the post-processing stage when computing Inv-Entropy scores. Furthermore, in settings where response generation time is negligible, such as when LLMs are deployed locally and perturbations are processed in parallel, our total computational cost is often lower than that of existing baselines.
Table 4: Computation efficiency comparison across methods.
Method Peak GPU Memory (MB) Time A (s) Time B (s)
Semantic Entropy 3575.31 13.985 20.183
VU 0 0 0.620
P(True) 0 0.001 2.201
LexSim 0 0.022 3.189
DegMat 1610.29 6.143 15.905
LUQ 1580.21 3.821 4.432
KLE 1608.52 1.323 6.725
Inv-Entropy (Ours) 86.65 1.990 10.323
NR-Inv-Entropy (Ours) 86.65 1.769 6.358
D.3 Additional experimental results
Table 5 presents the complete TSU results across all five datasets.
[…]
Method
TSU (0.3, 0.7)
TSU (0.7, 1.0)
TSU (1.0, 1.4)
TSU (0.3, 1.0)
TSU (0.3, 1.4)
TSU (0.7, 1.4)
TSU (0.3-1.0)
TSU (0.7-1.4)
TSU (0.3-1.4)
TriviaQA
Semantic Entropy 11.57 14.26 17.35 21.56 25.51 20.64 5.18 5.35 3.94
VU 22.45 27.55 38.78 25.51 41.84 42.86 0.00 4.92 0.00
P(True) 1.02 2.13 3.85 1.38 0.98 3.49 0.00 0.00 0.00
LexSim 22.39 22.95 46.94 30.61 53.36 53.54 9.18 12.38 8.16
DegMat 24.58 33.21 45.37 31.77 48.98 47.96 18.37 20.02 13.27
LUQ 20.41 31.08 48.06 33.67 52.34 50.00 14.78 27.63 10.20
KLE 7.14 17.35 13.45 4.57 1.28 6.42 2.79 1.31 0.00
Inv-Entropy 52.42 66.33 77.55 76.53 92.86 88.78 30.49 47.21 19.05
NI-Entropy 55.14 47.96 61.22 59.35 67.35 60.20 21.43 19.32 11.73
WD-px-py 44.81 67.35 57.62 60.20 64.49 69.39 23.11 36.73 12.06
MAX-py-x 65.31 50.83 74.49 67.35 85.71 81.63 27.42 32.65 16.03
SciQ
Semantic Entropy 11.09 15.48 21.55 17.35 27.62 25.51 6.73 4.31 2.65
VU 21.43 33.67 28.57 35.71 31.42 32.65 1.55 6.12 0.00
P(True) 0.00 3.42 3.14 0.00 1.33 0.79 0.00 0.00 0.00
LexSim 31.73 29.49 47.96 40.82 61.22 59.18 19.39 13.27 11.22
DegMat 39.80 47.96 57.14 52.04 64.28 64.46 23.47 23.99 13.04
LUQ 38.45 43.88 41.84 44.90 57.14 44.95 18.37 17.35 6.12
KLE 6.12 39.54 19.32 4.98 5.67 18.37 2.30 5.10 0.00
Inv-Entropy 46.70 62.24 72.38 66.72 75.84 80.61 22.45 35.71 14.31
NI-Entropy 42.30 59.18 55.19 48.98 54.77 63.27 21.43 22.65 9.18
WD-px-py 50.00 59.16 63.25 59.37 71.40 69.39 22.13 28.59 11.22
MAX-py-x 51.02 61.22 67.35 61.22 69.86 73.25 24.73 34.69 15.29
NQ
Semantic Entropy 22.53 20.41 35.71 27.55 39.80 30.78 4.08 10.20 2.04
VU 23.71 22.68 40.21 21.65 36.08 36.08 4.12 3.09 0.00
P(True) 4.08 3.06 2.04 1.02 2.04 1.02 0.00 0.00 0.00
LexSim 51.55 54.64 59.79 63.73 79.38 74.23 36.08 32.99 17.44
DegMat 51.04 51.04 59.38 65.62 76.04 71.88 32.29 33.33 21.42
LUQ 43.16 49.47 67.37 57.89 76.84 74.74 24.21 30.41 13.68
KLE 6.38 28.72 21.28 6.38 4.26 12.77 2.13 4.26 0.00
Inv-Entropy 61.86 59.79 78.35 74.23 91.75 85.57 30.93 41.24 22.43
NI-Entropy 47.42 60.82 59.79 58.76 70.10 70.10 23.71 29.81 10.31
WD-px-py 63.55 50.52 70.10 67.01 76.29 72.16 27.84 25.77 16.92
MAX-py-x 55.67 62.89 73.20 66.30 84.54 82.47 28.87 39.18 18.56
MMLU
Semantic Entropy 21.43 17.35 33.20 21.83 42.08 39.80 7.14 4.93 2.09
VU 23.56 19.01 37.62 30.18 32.50 38.73 1.37 2.59 0.00
P(True) 1.92 4.56 5.87 4.92 5.02 5.79 0.00 0.00 0.00
LexSim 33.94 41.17 55.06 50.00 68.37 61.22 24.78 30.61 15.28
DegMat 41.84 54.08 69.39 53.76 78.57 77.55 21.46 32.58 14.34
LUQ 53.46 47.96 61.22 58.92 68.37 62.24 27.55 27.55 10.80
KLE 10.20 25.51 26.53 7.14 4.76 12.23 2.93 2.67 0.00
Inv-Entropy 60.08 67.35 73.47 79.59 90.82 86.73 34.31 43.88 18.37
NI-Entropy 56.12 54.63 50.00 67.35 52.45 59.38 19.39 18.37 7.14
WD-px-py 50.00 59.16 66.38 59.37 70.41 69.39 25.60 21.65 11.22
MAX-py-x 62.24 62.35 65.41 73.47 81.73 80.61 31.62 25.51 14.42
GSM8K
Semantic Entropy 44.90 56.12 35.71 71.43 77.55 62.24 20.41 13.27 4.08
VU 11.34 39.18 29.90 35.05 35.05 38.14 2.06 6.19 1.03
P(True) 5.10 17.35 6.12 3.32 3.98 11.22 0.00 0.00 0.00
LexSim 54.17 63.54 54.17 65.62 64.58 63.54 30.31 23.96 10.42
DegMat 55.79 64.21 55.79 72.63 75.79 70.53 29.47 29.47 10.53
LUQ 64.95 74.23 72.16 83.51 89.69 81.44 44.33 54.64 37.11
KLE 17.02 35.11 27.66 4.23 3.72 23.40 0.00 3.19 0.00
Inv-Entropy 73.68 67.33 68.42 86.32 93.68 72.63 45.81 44.21 30.53
NI-Entropy 53.61 58.76 51.55 61.16 57.73 59.79 25.77 20.62 6.19
WD-px-py 54.08 53.06 68.37 55.10 64.29 60.20 20.41 27.43 7.14
MAX-py-x 71.88 65.62 62.50 89.58 91.67 77.08 42.71 37.50 28.12
Table 6 shows the performance across different answer lengths in Natural Question (NQ) [kwiatkowski2019natural] dataset. NQ dataset includes both short and long-form answers. We selected questions that contain only long-form answers. The reference answers in this subset range from 34 to 350 tokens in length. To analyze the sensitivity of different answer lengths, we further divided the dataset into three subsets based on the number of tokens in the reference answers: Short (< 80 tokens), Medium (80–120 tokens) and Long ($\geq$ 120 tokens). We evaluated both baseline methods and our proposed method on each of these subsets as well as the full sample set.
Our method achieves state-of-the-art performance on the full dataset, ranking first across all three metrics. We also observe some sensitivity to answer length: it shows a clear and substantial advantage on the short-answer subset, highlighting the effectiveness of the inverse-design mechanism when responses are more concise. On the medium-length subset, it continues to outperform all baselines, though with a smaller margin. On the long-answer subset, while not the top performer, our method remains competitive and yields reasonable results compared to strong baselines such as LUQ.
Intuitively, these results align with expectations under the inverse perspective: the shorter the answer, the more important it becomes to explore the diversity of the input space. Nevertheless, our approach consistently ranks among the top two methods even in the long-form QA setting. We end by noting that our probabilistic framework is also generic by nature and can accommodate a forward perspective or alternative uncertainty metrics beyond entropy computed over $P(X\mid Y)$. Exploring such metrics across different answer lengths is a promising direction we leave for future work.
Table 6: Comparison of AUROC, PRR, and Brier scores across Short, Medium, Long, and Full subsets on NQ. Bold indicates the best performer. Underline indicates the second-best.
Metric (↑/↓) Method Short Medium Long Full
AUROC (↑) Semantic Entropy 0.509 0.461 0.584 0.521
VU 0.531 0.495 0.508 0.533
P(True) 0.529 0.473 0.548 0.519
LexSim 0.624 0.438 0.555 0.518
DegMat 0.547 0.621 0.484 0.551
LUQ 0.662 0.508 0.612 0.627
KLE 0.265 0.456 0.445 0.410
Inv-Entropy (Ours) 0.794 0.634 0.589 0.661
PRR (↑) Semantic Entropy 0.420 0.584 0.507 0.505
VU 0.489 0.615 0.495 0.537
P(True) 0.427 0.584 0.478 0.502
LexSim 0.628 0.684 0.544 0.563
DegMat 0.543 0.682 0.548 0.549
LUQ 0.649 0.655 0.523 0.595
KLE 0.354 0.649 0.444 0.449
Inv-Entropy (Ours) 0.747 0.742 0.508 0.614
Brier (↓) Semantic Entropy 0.227 0.230 0.221 0.242
VU 0.209 0.218 0.216 0.223
P(True) 0.225 0.232 0.230 0.244
LexSim 0.169 0.199 0.207 0.225
DegMat 0.187 0.181 0.213 0.229
LUQ 0.164 0.209 0.179 0.208
KLE 0.228 0.203 0.225 0.244
Inv-Entropy (Ours) 0.125 0.175 0.193 0.201
@@ D.3 Additional experimental results
[…]
@@
[…]
Table 10: Comparison of AUROC, PRR, and Brier scores on the TriviaQA dataset under varying temperatures ($\mathbb{T}={0.3,\ 0.7,\ 1.0,\ 1.4}$). We use GPT-3.5-Turbo with ChatGPT-based paraphrasing and the SBERT-small (paraphrase-MiniLM-L6-v2) embedding function. Bold and underline indicate the best and second-best performers, respectively.
@@
[…]
Table 11: Comparison of AUROC, PRR, and Brier scores on the SciQ dataset under varying temperatures ($\mathbb{T}={0.3,\ 0.7,\ 1.0,\ 1.4}$). We use GPT-3.5-Turbo with ChatGPT-based paraphrasing and the SBERT-large (all-mpnet-base-v2) embedding function. Bold and underline indicate the best and second-best performers, respectively.